# Chapter 0 Hadron Production Processes

### v1

@@ 2 Double-Pion Production and Polarization
An important set of observables on hadron structure and production dynamics is obtained in polarization measurements, see, e.g., the discussion around Fig. 29. For single pion production, such measurements are done at practically all hadron facilities and have become routine work in most of the CC approaches.
The situation is different, however, for double-pion production. The probably only double-pion production experiment including polarization measurements and subsequent theoretical analysis was done some time ago by Alekseev et al. [202] at ITEP (Moscow) . Data were taken for the reaction $\pi^{-}+\vec{p}\to\pi^{+}+\pi^{-}+p$ on polarized targets and at beam momentum of $1.78~{}\mathrm{Gev/c}$. By means of the SPIN spectrometer, designed especially for measurements of polarization observables in reactions leading to charged two- and three-particle final states, the full set of 14 spin observables could be measured. The primary aim of the experiment was to study specifically pion-pion interactions in the mass region of the $\rho(770)$ meson, but aiming at narrowing down the mass of the - until today - heavily disputed mass (and width) of the iso-scalar scalar $\sigma/f_{0}(500)$-meson (see Fig. 7 and the PDG2024 listings [38]).
The authors of Ref. [202] could indeed extract from the data highly valuable information on the $\pi^{+}\pi^{-}$ S-wave scattering phase shift in a PWA by inclusion of the polarization observables. However, the data obtained in that single measurement were not sufficient to determine unambiguously the desired mass parameter. Still, that - until today - single case experiment showed the important gain in information on the dynamics and spectroscopy of a two-body sub-system within the three-hadron final state, being populated in double-pion production on the nucleon.
@@ 3 Double Pseudoscalar Meson Production Beyond Pions
Double-kaon or eta meson production as well as vector meson (photo-)production are experimentally and theoretically highly demanding tasks. In a recent note, the ALICE collaboration announced first data on pion- and kaon-pair production [203]. From their mass spectra, they concluded that the states, which were observed in the mass region of $2~{}\mathrm{GeV}$, are well described by CQM calculations in the tradition of Karl and Isgur.
There is also ongoing research activity on the theoretical side. In [204] the authors presented a new approach describing double pion photoproduction off the nucleon in covariant chiral perturbation theory, thus connecting two important concepts of hadron physics.
The recent BnGa paper reported a combined analysis of photo and pion-induced double pion production [87]. Recent CBELSA/TAPS and CLAS (photo case) and Crystal Ball at BNL and HADES at CERN (pion case) collaboration data plus $\pi N$ PWA amplitudes from the SAID and Karlsruhe-Helsinki groups were involved in the analysis. The critical motivation of this study is the separation of $\pi\Delta$, $\rho N$, and $f^{0}(500)N$ final states because the experimental state for each case is the same, namely two pions and a nucleon. The CLAS experiment reported recently double-pion electroproduction on the proton and on deuterium [205], thus expanding the research into the domain of finite momentum transfer.
@@ 11 Hadron Production Induced by Nucleons, Nuclei, and Lepton Beams
Past measurements involving pion and kaon scattering measurements were made at a variety of laboratories, mainly in the 1970s and 1980s when experimental techniques were far inferior to the standards of today. In the US, pion beams in the momentum range $190~{}\mathrm{MeV/c}$ to $730~{}\mathrm{MeV/c}$ were available at the “Meson Factory” LAMPF in Los Alamos. This means that the maximum c.m. energy for baryon spectroscopy measurements at LAMPF was only $W\approx 1500~{}\mathrm{MeV}$. LAMPF was a linear accelerator for $1000~{}\mathrm{\mu A}$ of protons at $800~{}\mathrm{MeV}$. The meson factory PSI (former SIN) near Zurich was a sector-focused cyclotron capable of $100~{}\mathrm{\mu A}$ of protons at $600~{}\mathrm{MeV}$, and the Meson Factory TRIUMF in Vancouver was a sector-focused cyclotron for negative hydrogen ions up to $100~{}\mathrm{\mu A}$ at $500~{}\mathrm{MeV}$ [206, 207].
@@
Important work is being done by experiments at other hadron facilities which were not mentioned in detail. Historically, SATURN at Saclay and the Synchro-Phasotron in Dubna, (JINR) for example, gave important insight into meson production processes on nuclei. The STAR@RHIC (BNL) and the GSI at Darmstadt experiments produced - and are producing - a wealth of data on meson and kaon, and partially also vector meson production in heavy ion collisions at relativistic energies. The meson yields of such processes, occurring in compressed and/or heated nuclear matter, were essential for understanding the dynamics of stressed baryon matter. Practically all currently operating LHC experiments are searching for exotic hadrons and nuclei, exploring, e.g., flavored nuclei and antimatter-nuclei, the latter used for searching signals of CP and CPT violation. COMPASS was and AMBER is devoted to meson physics, both located at CERN. J-PARC at Tokai has a strong program on hadron spectroscopy. An own class of facilities are the electron-positron colliders BESIII and BELLE. The experiments are mainly focused on flavor physics.
1 Resonance Spectrum of the Nucleon
Coupled channels approaches like the ones discussed in the previous sections are a valuable tool for understanding the spectrum of the excited nucleon. A natural question is how well the spectra of CC approaches agree with the results of other independent spectroscopic investigations?
Here, we address that issue for the case of the GiM coupled channels approach which may serve as a representative example. As mentioned before, the GiM is based on a phenomenological covariant field theory of baryons, mesons, and their interactions. By construction, the fundamental symmetries of QCD, including also chiral symmetry, are conserved. Resonance and background contributions are generated consistently out of the tree-level interactions as defined by the underlying Lagrangian.
The scattering amplitudes are determined by a linear system of coupled equations which is solved numerically in K-matrix approximation and partial wave representation. The gauge-invariant description of high-spin resonances was discussed in detail. Applications to selected reaction channels have been presented, ranging from single pion, eta, and kaon production to double-pion production to investigations of omega and $K^{*}$ vector meson channels.
Hence, CC approaches like the GiM and the other aforementioned projects incorporate the defining elements of hadron physics on the theoretical level. The approaches use comparable numerical methods for solving the CC problem where attention is paid to the fact that the numerical procedures conserve the basic symmetries and conservation laws.
The GiM baryon level scheme resulting from the investigations of meson production channels is summarized in Fig. 34 and compared to the observed spectrum as found in the PDG resonance compilations [182, 38]. It is noteworthy that the GiM spectrum is the result of a network of calculations in which resonances of one kind serve as intermediate states in analyses of resonances in reactions of another kind, e.g., the $\pi\Delta$ channels in double-pion production. The rather satisfactory agreement is a strong and encouraging confirmation of CC approaches as a tool for spectroscopic research in hadron production reactions.
@@ 1 Resonance Spectrum of the Nucleon
12 Model-Independent Analysis and Optimized Storage of Hadronic Reaction Data
@@
The previous section was devoted to an involved multi-channel description of experimental data. For that purpose, models about the reaction mechanism and the interactions had to be invented, mostly on phenomenological grounds. Hence, model parameters were adjusted by fits to data. As seen, those approaches are successful in the sense of reproducing data in a consistent manner which, in view of the complexity of the task, is a remarkable achievement.
@@
However, modern accelerator facilities in combination with detectors are capable of providing a tremendous amount of experimental data. Here the problem arises, how to present those numerous detailed data in a model-independent manner and make them available - and manageable - for further investigations. The standard approach is graphical representation, displaying various cuts through the multi-dimensional volume of measured data resulting in a multitude of graphs. That approach is necessarily selective and lacking complete coverage.
@@
As an alternative, Azimov et al. [208] has suggested expanding experimental data into a series of a complete set of orthogonal functions, e.g., Legendre polynomials $P_{J}(z)$. For unpolarized differential cross sections, measured angular distributions, for example, may be decomposed as
@@
where where $W$ is the c.m. energy, $z=\cos\theta$, and $\theta$ is the polar c.m. angle. Formally, this series is infinite. However, in real situations only a finite number of Legendre coefficients $A^{(\sigma)}_{J}(W)$ are effectively necessary. As a cut-off criterion one may impose that only those are considered with coefficients whose central values are larger by a given limit than the estimated error. The method was illustrated, for instance, in Ref. [209] using photoproduction data of the reaction $\gamma p\to\pi^{0}p$. As a result, the whole set of data appears to be represented in a compact form of energy dependent Legendre coefficients (Fig. 35), easy to use in later theoretical re-analyses or any other application of the data. The approach may be optimized to other kinds of data by choosing function systems well adapted to the case. Such compression strategies are used widely in modern data management and data mining.
@@
The Legendre coefficients have also other, less evident properties. For parity conserving reactions, some of those properties are related to partial-wave amplitudes for states of definite parity. In the unpolarized cross section the positive- and negative-parity amplitudes always appear symmetrically (that is why the unpolarized cross section by itself does not allow to determine the parity of a particular state). It is not quite so for $A^{(\sigma)}_{J}$. One can show that the Legendre coefficients provide specific discrimination of parities: $A^{(\sigma)}_{J}$ with odd $J$ contain only interferences of states with opposite parities, while $A^{(\sigma)}_{J}$ with even $J$ contain only interferences of states with the same parities, positive or negative. And, of course, only the even $J$ coefficients may contain squares of absolute values of various partial-wave amplitudes.
One can summarize the above statements in the following way:
  • Legendre expansions provide a model-independent approach suitable for presentation of modern detailed (high-precision and high-statistics) data for two-hadron reactions.
  • This approach is applicable both to cross sections and to polarization variables; it is much more compact than traditional methods, at least, at not very high energies.
@@
  • Due to interference with resonances, high momentum Legendre coefficients open a window to study higher partial-wave amplitudes, which are out-of-reach in any other ways.
Concluding this, one should emphasize that direct interference has become a useful instrument to search for and study rare decays of well-established resonances. However, its possibilities are limited by restrictions for the resonance quantum numbers. Rescattering interference is not limited by such requirements and, therefore, may provide effective methods to search for and study new resonances with arbitrary quantum numbers. Data on multi-hadron decays of heavy particles also present a new rapidly expanding area for applications of different kinds of interference both to study the spectroscopy of resonances and to establish their characteristics.
@@ 13 Summary and Outlook
The long journey from first, at that time not understood signals observed by Victor in high-altitude balloon campaigns, and in the following years increasingly being confirmed in different independent experiments to modern hadron physics was reviewed. The improving experimental techniques were accompanied by epochal developments in theory, changing forever our understanding of the material world and the origin of matter. That exciting journey is by far not coming to an end - it is ongoing and many surprises can be expected for the future.
The final goal - yet waiting to be achieved - is to connect the results derived by CC methods from data to QCD predictions as obtained by LQCD and the respective functional approaches discussed in this Encyclopedia. The iterative approach, eventually converging to a (self-) consistent picture, is indicated in Fig. 36.

### v2

@@ 3 Double-Pion Production and Polarization
An important set of observables on hadron structure and production dynamics is obtained from polarization measurements; see, e.g., the discussion around Fig. 21. For single-pion production, such measurements are conducted at practically all hadron facilities and have become routine work in most charged current (CC) approaches.
The situation is different, however, for double-pion production. The probably only double-pion production experiment, including polarization measurements and subsequent theoretical analysis, was conducted some time ago by Alekseev et al. [228] at ITEP (Moscow). Data were taken for the reaction $\pi^{-}\vec{p}\to\pi^{+}\pi^{-}+p$ on polarized targets and at a beam momentum of $1.78~\mathrm{Gev/c}$. By means of the SPIN spectrometer, designed especially for measurements of polarization observables in reactions leading to charged two- and three-particle final states, the full set of 14 spin observables could be measured. The primary aim of the experiment was to study specifically pion-pion interactions in the mass region of the $\rho(770)$ meson, while aiming to narrow down the mass of the - until today - heavily disputed mass (and width) of the iso-scalar scalar $\sigma/f_{0}(500)$-meson (see Fig. 7 and the listings in the 2024 issue of the Review of Particle Physics [41]).
The authors of Ref. [228] could indeed extract highly valuable information from the data on the $\pi^{+}\pi^{-}$ S-wave scattering phase shift in a PWA by including the polarization observables. However, the data obtained from that single measurement were not sufficient to determine the desired mass parameter unambiguously. Still, that - until today - single case experiment showed the important gain in information on the dynamics and spectroscopy of a two-body sub-system within the three-hadron final state, populated in double-pion production on the nucleon.
@@ 4 Double Pseudoscalar Meson Production Beyond Pions
Double-kaon or $\eta$-meson production, as well as vector meson (photo) production, is a task that is highly demanding, both experimentally and theoretically. In a recent note, the ALICE collaboration announced the first data on the pion and kaon pair production [229]. From their mass spectra, they concluded that the states observed in the mass region of $2~\mathrm{GeV}$ are well described by CQM calculations in the tradition of Karl and Isgur.
There is also ongoing research activity on the theoretical side. In [230], the authors presented a new approach describing double pion photoproduction off the nucleon in covariant chiral perturbation theory, thus connecting two important concepts of hadron physics.
The recent BnGa paper reported a combined analysis of photo- and pion-induced double pion production [99]. Recent CBELSA/TAPS and CLAS (photo case) and Crystal Ball at BNL and HADES at CERN (pion case) collaboration data, plus $\pi N$ PWA amplitudes from the SAID and Karlsruhe-Helsinki groups, were involved in the analysis. The critical motivation of this study is the separation of $\pi\Delta$, $\rho N$, and $f^{0}(500)N$ final states, because the experimental state for each case is the same: namely, two pions and a nucleon. The CLAS experiment recently reported double-pion electroproduction on the proton and on deuterium [231], thus expanding the research into the domain of finite momentum transfer.
5 Pion Production Spectra and Resonance Spectrum of the Nucleon
A natural and perhaps the most important question in hadron physics is how well the experimental spectrum is described by theory.
Here, we address this issue in the context of the GiM coupled channels approach, which may serve as a representative example for this class of models. As mentioned before, the GiM is based on a phenomenological covariant field theory of baryons, mesons, and their interactions. By construction, the fundamental symmetries of QCD, including chiral symmetry, are conserved. Resonance and background contributions are generated consistently from the tree-level interactions as defined by the underlying Lagrangian.
Figure 41: Comparison of GiM $C$-results (lines) to the data set of measured meson production spectra (symbols) observed in $\pi+p$ reactions.
The scattering amplitudes are determined by a linear system of coupled equations, which is solved numerically in the K-matrix approximation and partial wave representation. High-spin resonances are incorporated in a gauge-invariant manner. Applications to selected reaction channels have been presented, ranging from single pion, eta, and kaon production to double-pion production to investigations of omega and $K^{\ast}$ vector meson channels.
Hence, CC approaches like the GiM and the other aforementioned projects incorporate the defining elements of hadron physics at both the theoretical and numerical levels. The approaches use comparable numerical methods for solving the CC problem, where attention is paid to the fact that the numerical procedures conserve the basic symmetries and conservation laws.
A direct reliability test of CC models is the comparison of numerical results to data recording the spectra in the various channels under theoretical scrutiny. In Fig.41 a variety of total meson production cross sections observed in pion-proton reactions are compared to results of GiM CC calculations. Overall the large set of divers data is quite well described, hence confirming convincingly the validity of CC approaches like the GiM to hadron production.
The GiM baryon level scheme resulting from the investigations of meson production channels is summarized in Fig. 42 and compared to the observed spectrum, as found in the PDG resonance compilations [167, 41]. It is noteworthy that the GiM spectrum is the result of a network of calculations in which resonances of one kind serve as intermediate states in the analyses of resonances in reactions of another kind, e.g., the $\pi\Delta$ channels in double-pion production. The rather satisfactory agreement is a strong and encouraging confirmation of CC approaches as a tool for spectroscopic research in hadron production reactions.
@@ 10 Hadron Production Induced by Nucleons, Nuclei, and Lepton Beams
Past measurements involving pion and kaon scattering were made at a variety of laboratories, mainly in the 1970s and 1980s, when experimental techniques were far inferior to the standards of today. In the US, pion beams in the momentum range $190~\mathrm{MeV/c}$ to $730~\mathrm{MeV/c}$ were available at the “Meson Factory” LAMPF in Los Alamos. This means that the maximum c.m. energy for baryon spectroscopy measurements at LAMPF was only $W\approx 1500~\mathrm{MeV}$. LAMPF was a linear accelerator for $1000~\mathrm{\mu A}$ protons at $800~\mathrm{MeV}$. The meson factory PSI (formerly SIN) near Zurich was a sector-focused cyclotron capable of $100~\mathrm{\mu A}$ of protons at $600~\mathrm{MeV}$, and the Meson Factory TRIUMF in Vancouver was a sector-focused cyclotron for negative hydrogen ions up to $100~\mathrm{\mu A}$ at $500~\mathrm{MeV}$ [232, 233]. Single-pion production was also part of the WASA@COSY research program [234].
@@
Important work is being conducted through experiments at other hadron facilities that were not mentioned in detail. Historically, SATURN at Saclay and the Synchro-Phasotron in Dubna (JINR), for example, provided important insights into meson production processes on nuclei. The STAR@RHIC (BNL) experiment and the HADES@GSI experiment produced - and are producing - a wealth of data on meson and kaon, and partially also vector meson production in heavy ion collisions at relativistic energies. The meson yields of such processes, occurring in compressed and/or heated nuclear matter, are essential for understanding the dynamics of compressed and heated baryon matter.
Practically all currently operating LHC experiments are searching for exotic hadrons and nuclei, exploring, e.g., flavored hypernuclei and antimatter-nuclei, the latter used for searching signals of CP and CPT violation. COMPASS and AMBER Collaborations are devoted to meson physics; both are also located at CERN. J-PARC at Tokai has a strong program in hadron spectroscopy and strangeness studies [235].
A different class of facilities is the electron-positron colliders BESIII and Belle II, utilizing electro-weak interactions for hadron production in exit channel configurations of total $J^{\pi}=1^{-}$ as dictated by the producing annihilation reaction. The experiments are primarily focused on flavor physics, see e.g.[236, 237].
@@
11 Model-Independent Analysis and Optimized Treatment of Hadronic Reaction Data
@@
The previous section was devoted to an involved multi-channel description of experimental data. For that purpose, models of the reaction mechanism and the interactions had to be developed, mostly on phenomenological grounds. Hence, model parameters were adjusted by fitting them to the data. As seen, those approaches are successful in the sense of reproducing data in a consistent manner, which, in view of the complexity of the task, is a remarkable achievement.
@@
However, modern accelerator facilities, in combination with detectors, are capable of providing a tremendous amount of experimental data. Here, the problem arises: how to present those numerous detailed data in a model-independent manner and make them available - and manageable - for further investigations. The standard approach is a graphical representation, displaying various cuts through the multidimensional volume of measured data, resulting in a multitude of graphs. That approach is necessarily selective and lacks complete coverage.
@@
As an alternative, Azimov et al. [238] have suggested expanding experimental data into a series of a complete set of orthogonal functions, e.g., Legendre polynomials $P_{J}(z)$. For unpolarized differential cross sections, measured angular distributions, for example, may be decomposed as
@@
where $W$ is the c.m. energy, $z=\cos\theta$, and $\theta$ is the polar c.m. angle. Formally, this series is infinite. However, in real situations, only a finite number of Legendre coefficients $A^{(\sigma)}_{J}(W)$ are effectively necessary. As a cut-off criterion, one may impose that only those are considered with coefficients whose central values are larger by a given limit than the estimated error. The method was illustrated, for instance, in Ref. [239] using photoproduction data of the reaction $\gamma p\to\pi^{0}p$. As a result, the entire set of data appears to be represented in a compact form of energy-dependent form. Legendre coefficients (Fig. 43), easy to use in later theoretical re-analyses or any other application of the data. The approach may be optimized for other kinds of data by choosing function systems that are well adapted to the case. Such compression strategies are widely used in modern data management and data mining.
@@
The Legendre coefficients also have other, less evident properties. For parity-conserving reactions, some of these properties are related to partial-wave amplitudes for states of definite parity. In the unpolarized cross section, the positive- and negative-parity amplitudes always appear symmetrically (that is why the unpolarized cross section by itself does not allow for determining the parity of a particular state). It is not quite the same for $A^{(\sigma)}_{J}$. One can show that the Legendre coefficients provide specific discrimination of parities: $A^{(\sigma)}_{J}$ with odd $J$ contains only interferences of states with opposite parities, while $A^{(\sigma)}_{J}$ with even $J$ contains only interferences of states with the same parities, positive or negative. And, of course, only the even $J$ coefficients may contain squares of the absolute values of various partial-wave amplitudes.
The above statements can be summarized as follows:
  • Legendre expansions provide a model-independent approach suitable for the presentation of modern detailed (high-precision and high-statistics) data for two-hadron reactions.
  • This approach applies to both cross sections and polarization variables; it is much more compact than traditional methods, at least at not very high energies.
@@
  • Due to interference with resonances, high-momentum Legendre coefficients open a window to study higher partial-wave amplitudes, which are beyond reach in any other way.
In conclusion, it is worth emphasizing that direct interference has become a useful instrument for searching for and studying rare decays of well-established resonances. However, its possibilities are limited by restrictions on the resonance quantum numbers. Rescattering interference is not limited by such requirements and, therefore, may provide effective methods to search for and study new resonances with arbitrary quantum numbers. Data on multi-hadron decays of heavy particles also present a rapidly expanding area for applications of different kinds of interference, both to study the spectroscopy of resonances and to establish their characteristics.
A special kind of challenge is the treatment and analysis of large data sets, both from large scale experiments and numerical studies. Graph neural networks, especially designed and trained artificial intelligence (AI) systems, are indispensable tools for handling the formidable tasks of data evaluation, pattern recognition, and archiving. In [240], the status and perspectives of using graph neural networks, for example, at LHC experiments, are discussed as a flexible and efficient approach to encoding heterogeneous information. In [241], an initiative on AI infrastructure for particle, nuclear, and astroparticle physics was started recently. The authors strongly emphasize that Artificial Intelligence (AI) is transforming scientific research with deep learning methods. Applications will play a central role in data analysis, simulations, and signal detection across the fields of physics.
@@ 12 Summary and Outlook
The long journey from the first, at that time not understood, signals observed by Victor Hess in high-altitude balloon campaigns, which were increasingly confirmed in different independent experiments in the following years, to modern hadron physics was reviewed. The improving experimental techniques were accompanied by epochal developments in theory, changing our understanding of the material world and the origin of matter. That exciting journey is by no means coming to an end - it is ongoing, and many surprises can be expected in the future.
The final goal - yet waiting to be achieved - is to connect the results derived by CC methods from data to the QCD predictions obtained by LQCD and the respective functional approaches discussed in this Encyclopedia. The iterative approach, eventually converging to a (self-) consistent picture, is indicated in Fig. 44.